From GlueXWiki
Jump to: navigation, search

March 26, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Power outage issues
  2. Gas puzzle
  3. CDC update (Beni, Mike) CDC_recent_run_drift_times
  4. FDC update (Lubomir)
  5. Electronics (Fernando, Cody, Chris, Nick)
  6. Engineering
  7. Other


Participants: Naomi (CMU), Fernando, Dave, Chris, Nick, Cody, Mike, Beni, Simon, and Lubomir (JLab).

Power outage issues

- After restoring the FDC LV, many channels showed "temperature failure". It turned out the fans of the LV crate (U2-2-BOT) were not blowing. After unplugging/plugging the power cord, the fans started operating. Fernando: most likely it is a control issue, also one has to find a way to read the fan speed remotely. The CDC LV crate had no problems.

- After the outage, the chiller was turned on without touching the switch-off box, even the interlock should have prevented the box from turning on. Dave and Hovanes will look into this.

- All the elements of the gas system were of UPS and it never stopped for the FDC. Beni turned off the CDC MFCs in the gas room because they were in a continuous mode (temporary), therefore the pressure in the CDC mixing tank went to zero. It would be interesting to measrue the Oxygen contamination in the CDC.

Gas puzzle

- The new pump for the CO2 sensor will come tomorrow. After calibration the sensor will be installed in the Hall to monitor the adjustment of the CDC gas mixture to 50/50.

CDC update

- Naomi showed interesting results (linked above) comparing two runs before/after Beni made several modifications of the gas mixing system. They demonstrate change of the drift time in the CDC by ~10%, but most noticeably big change in the gain. By analyzing more runs in between we should figure out what caused the changes.

FDC update

- Lubomir studied the cathode resolution using the runs with 55Fe source, both with the FDC in the Hall and with the spare package. Cathodes can reconstruct the position of the wire (that was used as a trigger) with a surprisingly high precision. First plot shows the residual vs coordinate along the wire, where one see interesting structures with widths of ~100 microns. Even simple projection (second plot) gives a resolution of ~180 microns. These numbers should be divided by ~4 to get the cathode resolution along the wire, so we get 25-50 microns !?! Of course the advantage of the 55Fe source is that it gives a huge signal with a standard shape, which is not the case for the hits from the charged particle. Still this technique can be used to study systematic effects like cathode deformations.

- One HV channel (FDC4 cell 6 sector 2p) shows unstable HV, going down from time to time by up to 7-8Volts, and jumping current up to 2-3 uA (two plots attached). By swapping cables, it was found the problem is in the detector. The current can be explained by the detector, but it is not clear why the HV is not stable.


- Beni answering Cody's question: the flash 125s should stay ON when the DAQ is not running, they don't consume a lot of power.


- Yesterday morning we found a Fluorinert leak/drip at the connection of the return to the chiller. After fixing it one can still sniff leakage at almost all connections to the chiller, but since then the loss of Fluorinert was reduced significantly.